OpenStudy (anonymous):

locate the discontinuity y = 4/(9 + e^(7/x))

OpenStudy (watchmath):

x=0 only

OpenStudy (anonymous):

could you explain

OpenStudy (watchmath):

the function e^(7/x) is undefined at x=0, other than that e^(7/x) is continuous

OpenStudy (anonymous):

oh, so its discontinuous when the domain is undefined or zero?

OpenStudy (watchmath):

That is one of the case where the function discontinuous. Remember that the function is continuous at x=c if \(\lim_{x\to c}f(x)=f(c)\). So if we can't evaluate f(c), i.e., f(c) is undefined then that immediatly implies that the function is discontinuous at x=c.

OpenStudy (anonymous):

could you help with this _________discontinuity of y = ln((tan(x))2)